# Gongxu Luo1,  Yuewen Sun1,2,  Salman Khan1,4,  Peter Spirtes2,  Kun Zhang1,2

### v1

@@ Acknowledgements
Bundesliga Terms of Use Services $\rightarrow$ 8. Audiovisual Content (bun, ) “The audiovisual content available within the Products is made available to the User for personal and non-commercial purposes only. The User is authorized to use this audiovisual content only for the purposes of information and entertainment in the private sphere for themselves and persons personally associated with them (e.g. family members, friends and acquaintances). Limited to these purposes, the DFL grants the User a non-exclusive, non-transferable, non-sub-licensable right of use to access and view the audiovisual content within the Products. With the exception of the aforementioned limited right of use, the User is not granted any rights to the audiovisual content.” No
@@
Ligue 1 Terms and Conditions of Use $\rightarrow$ 6. Intellectual Property (lig, ) “… Any total or partial reproduction of the Site or its elements without prior written authorization from the publisher (LFP) may lead to legal proceedings against the infringers.” Yes
@@
PGA Tour Terms of Use $\rightarrow$ 7. Conduct(D) (pga, ) “You may use real time scoring, statistics and other data (whether current or archival) collected from PGATOUR.COM solely for legitimate news reporting and for personal, non-commercial purposes. You shall not use real time scoring, statistics or other data (whether current or archival) collected from PGATOUR.COM for sale, license or other commercial purposes (including, without limitation, commercial gambling purposes), unless expressly licensed by the PGA TOUR Parties.” No
@@
ATP Tour Terms & Conditions $\rightarrow$ 7. PROHIBITED USES $\rightarrow$ A. Ownership (atp, ) “ATP owns or has the right to use all of the data, information, text, images, streaming media, video, sounds, icons, scores, rankings, statistics, and other content contained on this Website (the “Content”), and the copyrights and other intellectual property rights therein, unless otherwise noted. You may print one copy of the Content of this Website for your own personal, non-commercial use.” No
@@
The Figure LABEL:app_ahtle_distribution shows the distribution of AthlePersona. The AthlePersona dataset is dominated by athletes born between 1985 and 2005, with a uniform spread across birth months and days. Most individuals are associated with NFL, NHL and NBA, showing a strong skew toward U.S. sports. Heights cluster around 180–199 cm, and weights around 90–109 kg, which aligns with physical norms for elite athletes in contact sports. Nationalities are overwhelmingly North American, with minimal representation from other continents, highlighting a clear Western and U.S.-centric dataset bias.
The Figure LABEL:app_celeb_distribution shows the distribution of CelebPersona. The CelebPersona dataset predominantly features younger individuals, with birth years peaking between 1990–1999, and shows a balanced distribution across birth months and days. Most individuals are from North America and Europe, with underrepresentation from other continents. There is a notable occupational bias toward Entertainment, Music, and Sports, while fields like Healthcare and Academia are sparsely represented. Females slightly outnumber males. Height and weight distributions center around typical adult ranges, though outliers exist. The weight distribution peaks between 60–69 kg and 50–59 kg, with a sharp drop after 90 kg. The range 135–139 kg and higher has a minimal count. Regarding facial attributes, the majority of features are marked as absent, with a smaller subset present, particularly for traits like Oval Face and High Cheekbones. The unknown values are minimal.
@@
Throughout this paper, we rely on large language models (LLMs) to generate human personality. Fortunately, Benefiting from the recent explosion in the size and availability of LLMs (Achiam et al., 2023; Floridi & Chiriatti, 2020; Jiang et al., 2024; Liu et al., 2024; Bai et al., 2023), some research has shown that personality measurements in the outputs of some LLMs under specific prompting configurations are valid and reliable (Serapio-Garc´ıa et al., 2023; Jiang et al., 2023; Tseng et al., 2024). We built our datasets from a number of professional athletes and celebrities, based on the facts that they are famous and it is likely to have sufficient information about them online.
@@
We present the comparative evaluations on 10 of the state-of-the-art LLMs on our two persona datasets. Initially, we choose the Top 10 models based on the Arena leaderboard (Chiang et al., 2024). To enhance diversity, we also included Qwen2.5-Max and QwQ-32B (Bai et al., 2023) from Alibaba, both noted for their strong reasoning capabilities. We list all those 12 LLMs and summarize them in Table A5, including the model name, the company name, the arena score, the API input and output price per million tokens, and whether it is used by us for further analysis in Table 1. Specifically, GPT-4.5-Preview (Achiam et al., 2023) was excluded due to prohibitively high API costs where the input and output API prices were \$75 and \$150 per million tokens, respectively. Gemini-2.0-Pro-Exp-02-05 (Team et al., 2023) was omitted due to inaccessibility, it was merged to the latest Gemini-2.5-Pro-Exp-03-25 model. Therefore, in the end, we only considered 10 LLMs, as shown in Table 1.
@@
Prior research has demonstrated the importance of prompt engineering strategies in enhancing LLM performance across various tasks (Serapio-Garc´ıa et al., 2023; Jiang et al., 2023; Tseng et al., 2024; Wang et al., 2023; Li et al., 2024a; Tang et al., 2025). These foundational studies have established that prompt structure and presentation significantly influence model outputs, particularly in psychological assessment applications. Building on this foundation, we systematically investigate how variations in scoring scale format affect the consistency of LLM-generated trait assessments across the Big Five traits, with implications for reliable automated psychological evaluation.
@@
To evaluate independence relationships across different variable types in our analysis, we employ five statistical testing methods. For discrete variables, we utilize two classical approaches: the Chi-square test (Tallarida et al., 1987), which evaluates statistical independence between categorical variables, and the G-square test (Tsamardinos et al., 2006), a likelihood-ratio variant that demonstrates improved robustness in small sample scenarios. For continuous and mixed variable types, we implement three kernel-based methods: the Hilbert-Schmidt Independence Criterion (HSIC) (Gretton et al., 2005), which measures dependence in high-dimensional data using reproducing kernel Hilbert spaces; the Randomized Conditional Independence Test (RCIT) (Strobl et al., 2019), a non-parametric approach that employs randomized Fourier features to approximate kernel-based conditional independence testing; and the Kernel-based Conditional Independence Test (KCI) (Zhang et al., 2012), which extends HSIC methodology for testing conditional independence in complex data structures. This comprehensive suite of methods enables robust independence testing across diverse data types encountered in our experimental framework. A dependency is deemed significant if $p<0.05$, and each cell in Fig. 3(a)/(b) shows the number of methods that detect such significance, and we summarize these 5 methods in Table LABEL:app_athlete_cit and Table LABEL:app_celeb_cit.
@@
Figure LABEL:app_athlete_cit presents heatmaps of p-values from different statistical independence. The Chi-Square Test (CSQ) and G-Square Test (GSQ) show remarkably similar patterns, which is expected given their shared theoretical foundation for categorical variables. Overall, the independence test analysis reveals limited but significant demographic-trait dependencies in the AthlePersona dataset. Most relationships show p-values well above the 0.05 significance threshold, indicating statistical independence between demographic features and behavior traits. However, notable exceptions include birth year and league’s strong dependence with all big five traits, birth month associations with Agreeableness in the CSQ test (p = 0.001), which represents the strongest dependency detected. Birth day shows somewhat clear independence with trait in most methods. The kernel-based methods (RCIT, HSIC, KCI) generally produce lower p-values, indicating stronger evidence for dependencies. Most relationships show p-values between 0-0.01, suggesting statistical dependence among variables such as birth year, league, latitude, longitude, weight, and the Big Five behavior traits. Interestingly, weight is more dependent on openness, agreeableness, and neuroticism, while being more independent of conscientiousness and extraversion.
@@
Figure LABEL:app_celeb_cit shows heatmaps of p-values from different statistical independence tests evaluating the relationship between facial/demographic features and Big Five personality traits in the CelebPersona dataset. Features like birth year, gender, occupation, latitude, longitude, pointy nose and big lips frequently show strong associations with Big Five traits. In contrast, attributes like birth day, narrow eyes and bushy eyebrows generally appear independent of traits.
@@
We first introduce another operator to represent the point-wise distributional transformation. To maintain generality, we denote two variables as $a$ and $b$, with respective support sets $\mathcal{A}$ and $\mathcal{B}$.
@@
(Linear Operator) (Dunford & Schwartz, 1971) Consider two random variables $a$ and $b$ with support $\mathcal{A}$ and $\mathcal{B}$, the linear operator $L_{b|a}$ is defined as a mapping from a probability function $p_{a}$ in some function space $\mathcal{F}(\mathcal{A})$ onto the probability function $p_{b}=L_{b|a}\circ p_{a}$ in some function space $\mathcal{F}(\mathcal{B})$,
@@
For brevity, we define $\mathbf{w}_{m}\coloneqq[\mathbf{z}_{m},\mathbf{s}]$ with the support set $\mathcal{W}_{m}$.
@@
$\mathbf{x}_{m,A},\mathbf{x}_{m,B},\mathbf{x}_{m,C}$ are conditional independent given $\mathbf{w}_{m}$, which implies two equations:
@@
We can obtain $p(\mathbf{x}_{m,C},\mathbf{x}_{m,B}\mid\mathbf{x}_{m,A})$ directly from the observations, $p(\mathbf{x}_{m,A})$ and $p(\mathbf{x}_{m,C},\mathbf{x}_{m,B},\mathbf{x}_{m,A})$, and then the transformation in density function are established by
@@
We begin by marginalizing out the variable $\mathbf{x}_{m,A}$ using the transformation structure defined in Equation 8:
@@
This joint density expression can be rewritten using linear operators as defined in Definition 1 and Definition 2:
@@
We now integrate both sides over $\mathbf{x}_{m,B}\in\mathcal{X}_{m,B}$:
@@
Since integrating out $\mathbf{x}_{m,B}$ amounts to marginalizing over the joint representation, we obtain:
@@
Assuming $L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}$ is injective (see Assumption ii), we can invert it and obtain:
@@
Substituting Equation 17 into the earlier composition in Equation 14, we derive:
$L_{\mathbf{x}_{m,B};\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}=L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}L^{-1}{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}L_{\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}.$ (18)
Multiplying both sides of Equation 18 by $L^{-1}{\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}$ yields:
$L_{\mathbf{x}_{m,B};\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}L^{-1}{\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}=L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}L^{-1}{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}.$ (19)
The R.H.S. of Equation 19 is in a canonical conjugation form. Under Assumption i and by the uniqueness of spectral decomposition (see (Conway, 1994), Ch. VII and (Dunford & Schwartz, 1971), Theorem XV.4.5), we have:
$L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}L^{-1}{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}$ $=$ (20)
$(CL_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}P)($ $P^{-1}D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}P)(P^{-1}L^{-1}{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}C^{-1}),$
where $C$ is a nonzero scalar and $P$ is an invertible operator encoding permutation of the eigenbasis.
This yields the following identification up to permutation and rescaling:
$L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}=CL_{\mathbf{x}_{m,C}\mid\hat{\mathbf{w}}_{m}}P,\quad D{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}=P^{-1}D_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}P.$ (21)
We obtain a unique spectral decomposition in  Equation 21 with permutation and scaling indeterminacies. In the following, we will show how these indeterminacies can be resolved—if not, what informative results can still be inferred.
Since the normalizing condition
@@
must hold for every $\hat{\mathbf{w}}_{m}$, one only solution of $\int_{\mathcal{X}_{m,C}}Cp_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}\,d\mathbf{x}_{m,C}=1$ is to set $C=1$.
After that, we start from $D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}=P^{-1}D_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}P$. The operator, $D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}$, corresponding to the set $\{p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m})\}$ for fixed $\mathbf{x}_{m,B}$ and all $\mathbf{w}_{m}$, admits a unique solution ($P$ only change the entry position):
$\{p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m})\}=\{p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m})\},\quad\text{ for all }\mathbf{w}_{m},\hat{\mathbf{w}}_{m}.$ (23)
Due to the set is unorder, the only way to match the R.H.S. with the L.H.S. in a consistent order is to exchange the conditioning variables, that is,
$\{p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m}^{(1)}),p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m}^{(2)}),\ldots\}$ $=$ (24)
$\{p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}^{(1)}),$ $p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}^{(2)}),\ldots\}$ (25)
$\implies$ $[p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m}^{(\pi(1))}),p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m}^{(\pi(2))}),\ldots]$ $=$ (26)
$[p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}^{(\pi(1))}),$ $p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}^{(\pi(2))}),\ldots]$ (27)
where superscript $(\cdot)$ denotes the index of a conditioning variable, and $\pi$ is reindexing the conditioning variables. We use a relabeling map $h$ to represent its corresponding value mapping:
$p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid h(\mathbf{w}_{m}))=p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}),\quad\text{ for all }\mathbf{w}_{m},\hat{\mathbf{w}}_{m}.$ (28)
By Assumption iii, different $\mathbf{w}_{m}$ corresponds to different $p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m})$, there is no repeated element in $\{p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m})\}$ (and $\{p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m})\}$). Hence, the relabelling map $h$ is one-to-one (invertible).
Furthermore, Assumption 4 implies that $p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid h(\mathbf{w}_{m}))$ determines a unique $h(\mathbf{w}_{m})$. The same holds for the $p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m})$, implying that
$p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid h(\mathbf{w}_{m}))=p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m})\implies\hat{\mathbf{w}}_{m}=h(\mathbf{w}_{m}).$ (29)

### v2

@@ Acknowledgements
Bundesliga Terms of Use Services $\rightarrow$ 8. Audiovisual Content (11) “The audiovisual content available within the Products is made available to the User for personal and non-commercial purposes only. The User is authorized to use this audiovisual content only for the purposes of information and entertainment in the private sphere for themselves and persons personally associated with them (e.g. family members, friends and acquaintances). Limited to these purposes, the DFL grants the User a non-exclusive, non-transferable, non-sub-licensable right of use to access and view the audiovisual content within the Products. With the exception of the aforementioned limited right of use, the User is not granted any rights to the audiovisual content.” No
@@
Ligue 1 Terms and Conditions of Use $\rightarrow$ 6. Intellectual Property (74) “… Any total or partial reproduction of the Site or its elements without prior written authorization from the publisher (LFP) may lead to legal proceedings against the infringers.” Yes
@@
PGA Tour Terms of Use $\rightarrow$ 7. Conduct(D) (101) “You may use real time scoring, statistics and other data (whether current or archival) collected from PGATOUR.COM solely for legitimate news reporting and for personal, non-commercial purposes. You shall not use real time scoring, statistics or other data (whether current or archival) collected from PGATOUR.COM for sale, license or other commercial purposes (including, without limitation, commercial gambling purposes), unless expressly licensed by the PGA TOUR Parties.” No
@@
ATP Tour Terms & Conditions $\rightarrow$ 7. PROHIBITED USES $\rightarrow$ A. Ownership (7) “ATP owns or has the right to use all of the data, information, text, images, streaming media, video, sounds, icons, scores, rankings, statistics, and other content contained on this Website (the “Content”), and the copyrights and other intellectual property rights therein, unless otherwise noted. You may print one copy of the Content of this Website for your own personal, non-commercial use.” No
@@
The Figure 7(a) shows the distribution of AthlePersona. The AthlePersona dataset is dominated by athletes born between 1985 and 2005, with a uniform spread across birth months and days. Most individuals are associated with NFL, NHL and NBA, showing a strong skew toward U.S. sports. Heights cluster around 180–199 cm, and weights around 90–109 kg, which aligns with physical norms for elite athletes in contact sports. Nationalities are overwhelmingly North American, with minimal representation from other continents, highlighting a clear Western and U.S.-centric dataset bias.
The Figure 7(b) shows the distribution of CelebPersona. The CelebPersona dataset predominantly features younger individuals, with birth years peaking between 1990–1999, and shows a balanced distribution across birth months and days. Most individuals are from North America and Europe, with underrepresentation from other continents. There is a notable occupational bias toward Entertainment, Music, and Sports, while fields like Healthcare and Academia are sparsely represented. Females slightly outnumber males. Height and weight distributions center around typical adult ranges, though outliers exist. The weight distribution peaks between 60–69 kg and 50–59 kg, with a sharp drop after 90 kg. The range 135–139 kg and higher has a minimal count. Regarding facial attributes, the majority of features are marked as absent, with a smaller subset present, particularly for traits like Oval Face and High Cheekbones. The unknown values are minimal.
(b) The distributions of the features in CelebPersona Dataset.
@@
Throughout this paper, we rely on large language models (LLMs) to generate human personality. Fortunately, Benefiting from the recent explosion in the size and availability of LLMs (Achiam et al., 2023; Floridi and Chiriatti, 2020; Jiang et al., 2024; Liu et al., 2024; Bai et al., 2023), some research has shown that personality measurements in the outputs of some LLMs under specific prompting configurations are valid and reliable (Serapio-García et al., 2023; Jiang et al., 2023; Tseng et al., 2024). We built our datasets from a number of professional athletes and celebrities, based on the facts that they are famous and it is likely to have sufficient information about them online.
@@
We present the comparative evaluations on 10 of the state-of-the-art LLMs on our two persona datasets. The 10 LLMs include: Gemini-2.5-Pro-Exp-03-25, Llama-4-Maverick-03-26-Experimental, ChatGPT-4o-latest (2025-03-26), Grok-3-Preview-02-24, Gemini-2.0-Flash-Thinking-Exp-01-21, DeepSeek-V3-0324, DeepSeek-R1, Gemini-2.0-Flash-001, Qwen2.5-Max, and QwQ-32B. Initially, we choose the Top 10 models based on the Arena leaderboard (Chiang et al., 2024). To enhance diversity, we also included Qwen2.5-Max and QwQ-32B (Bai et al., 2023) from Alibaba, both noted for their strong reasoning capabilities. We list all those 12 LLMs and summarize them in Table A5, including the model name, the company name, the arena score, the API input and output price per million tokens, and whether it is used by us for further analysis in Table 1. Specifically, GPT-4.5-Preview (Achiam et al., 2023) was excluded due to prohibitively high API costs where the input and output API prices were \$75 and \$150 per million tokens, respectively. Gemini-2.0-Pro-Exp-02-05 (Team et al., 2023) was omitted due to inaccessibility, it was merged to the latest Gemini-2.5-Pro-Exp-03-25 model. Therefore, in the end, we only considered 10 LLMs, as shown in Table 1.
@@
Prior research has demonstrated the importance of prompt engineering strategies in enhancing LLM performance across various tasks (Serapio-García et al., 2023; Jiang et al., 2023; Tseng et al., 2024; Wang et al., 2023; Li et al., 2024a; Tang et al., 2025). These foundational studies have established that prompt structure and presentation significantly influence model outputs, particularly in psychological assessment applications. Building on this foundation, we systematically investigate how variations in scoring scale format affect the consistency of LLM-generated trait assessments across the Big Five traits, with implications for reliable automated psychological evaluation.
@@
(a) AthlePersona: Heatmap of P-value obtained from different independence test.
@@
To evaluate independence relationships across different variable types in our analysis, we employ five statistical testing methods. For discrete variables, we utilize two classical approaches: the Chi-square test (Tallarida et al., 1987), which evaluates statistical independence between categorical variables, and the G-square test (Tsamardinos et al., 2006), a likelihood-ratio variant that demonstrates improved robustness in small sample scenarios. For continuous and mixed variable types, we implement three kernel-based methods: the Hilbert-Schmidt Independence Criterion (HSIC) (Gretton et al., 2005), which measures dependence in high-dimensional data using reproducing kernel Hilbert spaces; the Randomized Conditional Independence Test (RCIT) (Strobl et al., 2019), a non-parametric approach that employs randomized Fourier features to approximate kernel-based conditional independence testing; and the Kernel-based Conditional Independence Test (KCI) (Zhang et al., 2012), which extends HSIC methodology for testing conditional independence in complex data structures. This comprehensive suite of methods enables robust independence testing across diverse data types encountered in our experimental framework. A dependency is deemed significant if $p<0.05$, and each cell in Fig. 3(a)/(b) shows the number of methods that detect such significance, and we summarize these 5 methods in Table 13(a) and Table 13(b).
(b) CelebPersona: Heatmap of P-value obtained from different independence test.
@@
Figure 13(a) presents heatmaps of p-values from different statistical independence. The Chi-Square Test (CSQ) and G-Square Test (GSQ) show remarkably similar patterns, which is expected given their shared theoretical foundation for categorical variables. Overall, the independence test analysis reveals limited but significant demographic-trait dependencies in the AthlePersona dataset. Most relationships show p-values well above the 0.05 significance threshold, indicating statistical independence between demographic features and behavior traits. However, notable exceptions include birth year and league’s strong dependence with all big five traits, birth month associations with Agreeableness in the CSQ test (p = 0.001), which represents the strongest dependency detected. Birth day shows somewhat clear independence with trait in most methods. The kernel-based methods (RCIT, HSIC, KCI) generally produce lower p-values, indicating stronger evidence for dependencies. Most relationships show p-values between 0-0.01, suggesting statistical dependence among variables such as birth year, league, latitude, longitude, weight, and the Big Five behavior traits. Interestingly, weight is more dependent on openness, agreeableness, and neuroticism, while being more independent of conscientiousness and extraversion.
@@
Figure 13(b) shows heatmaps of p-values from different statistical independence tests evaluating the relationship between facial/demographic features and Big Five personality traits in the CelebPersona dataset. Features like birth year, gender, occupation, latitude, longitude, pointy nose and big lips frequently show strong associations with Big Five traits. In contrast, attributes like birth day, narrow eyes and bushy eyebrows generally appear independent of traits.
@@
Figure A13: The high-level flowchart of the our theorems.
We first introduce an additional operator to represent pointwise distributional transformations, a concept widely used in the nonparametric identification literature (Hu and Schennach, 2008; Fu et al., 2025; Li et al., 2025b). To preserve generality, we denote any two variables by $a$ and $b$, with corresponding support sets $\mathcal{A}$ and $\mathcal{B}$, respectively.
@@
(Linear Operator) (Dunford and Schwartz, 1971) Consider two random variables $a$ and $b$ with support $\mathcal{A}$ and $\mathcal{B}$, the linear operator $L_{b|a}$ is defined as a mapping from a probability function $p_{a}$ in some function space $\mathcal{F}(\mathcal{A})$ onto the probability function $p_{b}=L_{b|a}\circ p_{a}$ in some function space $\mathcal{F}(\mathcal{B})$,
@@
For notational convenience, we define $\mathbf{w}_{m}\coloneqq[\mathbf{z}_{m},\mathbf{s}]$, with support set $\mathcal{W}_{m}$.
@@
The vectors $\mathbf{x}_{m,A}$, $\mathbf{x}_{m,B}$, and $\mathbf{x}_{m,C}$ are conditionally independent given $\mathbf{w}_{m}$, which implies the following two equations:
@@
We can directly obtain $p(\mathbf{x}_{m,C},\mathbf{x}_{m,B}\mid\mathbf{x}_{m,A})$ from the observed quantities $p(\mathbf{x}_{m,A})$ and $p(\mathbf{x}_{m,C},\mathbf{x}_{m,B},\mathbf{x}_{m,A})$. The corresponding density transformation is then given by
@@
We begin by marginalizing over the variable $\mathbf{x}_{m,A}$ using the transformation structure defined in Equation 8:
@@
This joint density can be equivalently expressed in terms of the linear operators defined in Definition 1 and Definition 2:
@@
Then, we integrate both sides over the $\mathbf{x}_{m,B}\in\mathcal{X}_{m,B}$:
@@
Since integrating out $\mathbf{x}_{m,B}$ corresponds to marginalizing over the joint representation, we obtain
@@
Assuming that $L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}$ is injective (see Assumption ii), we may invert this operator to obtain
@@
Substituting Equation 17 into the operator composition in Equation 14 yields
$L_{\mathbf{x}_{m,B};\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}=L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}\,D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}\,L^{-1}_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}\,L_{\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}.$ (18)
Multiplying both sides of Equation 18 by $L^{-1}_{\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}$ gives
$L_{\mathbf{x}_{m,B};\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}L^{-1}_{\mathbf{x}_{m,C}\mid\mathbf{x}_{m,A}}=L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}L^{-1}_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}.$ (19)
The right-hand side of Equation 19 takes a canonical conjugation form. Under Assumption i and by the uniqueness of spectral decomposition (see (Conway, 1994), Chapter VII, and (Dunford and Schwartz, 1971), Theorem XV.4.5), we have
$L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}L^{-1}_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}=(CL_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}P)(P^{-1}D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}P)(P^{-1}L^{-1}_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}C^{-1}),$ (20)
where $C$ is a nonzero scalar and $P$ is an invertible operator representing a permutation of the eigenbasis.
This yields identification up to permutation and scaling:
$L_{\mathbf{x}_{m,C}\mid\mathbf{w}_{m}}=CL_{\mathbf{x}_{m,C}\mid\hat{\mathbf{w}}_{m}}P,\qquad D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}=P^{-1}D_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}P.$ (21)
Equation Equation 21 provides a unique spectral decomposition up to permutation and scaling indeterminacies. We next show how these indeterminacies can be resolved, and when they cannot, what informative conclusions may still be drawn.
First, the normalization condition
@@
must hold for every $\hat{\mathbf{w}}_{m}$. Hence, the only solution is $C=1$.
Next, consider
$D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}=P^{-1}D_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}P.$
For fixed $\mathbf{x}_{m,B}$, the operator $D_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}$ corresponds to the collection $\{p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m})\}$ over all $\mathbf{w}_{m}$. Since $P$ only permutes entries, this collection admits a unique solution:
$\{p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m})\}=\{p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m})\},\quad\forall\,\mathbf{w}_{m},\hat{\mathbf{w}}_{m}.$ (23)
Because these sets are unordered, consistent matching requires a reindexing of the conditioning variables. Specifically,
$\{p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m}^{(1)}),\ldots\}$ $=\{p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}^{(1)}),\ldots\},$ (24)
$\implies[p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid\mathbf{w}_{m}^{(\pi(1))}),\ldots]$ $=[p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}^{(\pi(1))}),\ldots],$ (25)
where $\pi$ denotes a permutation of indices.
Let $h$ denote the corresponding relabeling map. Then
$p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid h(\mathbf{w}_{m}))=p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}),\quad\forall\,\mathbf{w}_{m},\hat{\mathbf{w}}_{m}.$ (26)
By Assumption iii, distinct $\mathbf{w}_{m}$ correspond to distinct conditional densities, implying that $h$ is one-to-one and invertible. Moreover, Assumption iii ensures that each conditional density uniquely determines its conditioning variable, so that
$p_{\mathbf{x}_{m,B}\mid\mathbf{w}_{m}}(\mathbf{x}_{m,B}\mid h(\mathbf{w}_{m}))=p_{\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m}}(\mathbf{x}_{m,B}\mid\hat{\mathbf{w}}_{m})\;\Longrightarrow\;\hat{\mathbf{w}}_{m}=h(\mathbf{w}_{m}).$ (27)